# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ 6 GBM eQTL Reanalysis
Four SNPs are identified by RegGMM to affect the network of co-expressions. Out of the four, rs1267622 is also identified by cspine to have a nonzero effect. It is interesting to look at the estimated effect of rs1267622, a variant of the BRAF gene, by the two methods, shown in Figure 4.
Figure 4: Estimated covariate networks for the SNP rs1267622, a variant local to the BRAF gene. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (3) and (7). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of GADD45A and GADD45B is detected by cspine while RegGMM detects an effect between PDGFRA and SOS2. The remaining two edges, between PDGRFA and PDGFB and PDGRFA and SOS2, are detected by both methods.
The results from both methods suggest that this SNP may modify the co-expressions of PDGFRA and PDGFB as well as PDGFRA and SOS2. This is plausible since all three of these genes lie in the Ras-Raf-MEK-ERK pathway along with BRAF (Table S3). There is evidence that variants on this pathway are associated with the proliferation of certain cancers [Gonzalez-Hormazabal et al., 2019]. Beyond this, the two methods differ; cspine estimates this variant to modify the co-expressions of GADD45A and GADD45B while RegGMM suggests that the co-expressions of PDGFRA and CAMK1 are modified.
[…]
The estimated connections among these four genes in the population network are shown in Figure 5. While RegGMM and cspine detect mostly the same edges in the population, only cspine finds that these co-expressions are modified by the presence of SNPs. These effects are summarized in Figure 5 (right). For instance, it is estimated that a variant local to the CALML4 gene may mediate the induction of the PI3K/Akt/mTOR pathway by expressions in E2F1. PIK3CA and E2F1 are part of interconnected signaling pathways implicated in cancer progression; E2F1 is involved in regulating the expression of PIK3CA and downstream signaling components [Ladu et al., 2008]. Hence, this discovery may give a clue for the role of CALML4 on regulating this pathway.
Figure 5: Estimated population network from cspine (left) and RegGMM (middle) highlighting four select genes. An edge indicates the partial correlation between the expressions of two genes conditional on the rest of the network (not shown). Right: the estimated effect of SNPs on co-expressions among the selected genes according to cspine. The gene to which the SNP is local is given in parentheses. Solid blue lines and dashed red lines indicate positive and negative edges, respectively. RegGMM did not detected any effects of SNPs on these co-expressions.
@@ 7 Discussion
In this work, we contributed to the covariate-adjusted Gaussian graphical model literature by developing a framework allowing for jointly convex optimization of the mean and precision matrix. Our theoretical work implies that the convex formulation allows for more relaxed scaling assumptions in the sparsities of $\boldsymbol{\mathbf{\beta}}_{j}$ in relation to the sample size $n$ and this is confirmed by our simulation results. Our method relies on tuning the pair $(\alpha_{s},\lambda_{0})$ over a grid as described in 3.3, which may be too computationally intensive for large datasets. If the practitioner has preconceptions about how many covariates are relevant in explaining the network structure of the responses, he may fix $\alpha_{s}$ to specify a desired group sparsity level. Alternatively, adapting tuning-free methods such as the square-root lasso [Belloni et al., 2011] to our method would be an important contribution for large datasets. A direction for theoretical work would be to establish the minimax rate of (15). Although our estimator achieves the same rate as in Zhang and Li [2022], the optimality of this rate has yet to be established. Our R package cspine is available on GitHub [^1] https://github.com/roobnloo/cspine.
@@ Acknowledgements
A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization
@@
[…]
       Define $\boldsymbol{\mathbf{W}}_{-j}$ as in (10) and solve
@@
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\}}+\tilde{\beta}_{kjh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\}},\quad j\neq k$
@@
Table S3 displays the genes belonging to known signaling pathways associated with GBM. The table is copied from Zhang and Li [2022]. Table S4 lists the SNPs that are detected by cspine to affect network co-expression.
@@
CALM1,CALML3, CALML4, CALML5, Maklad et al. [2019]
@@
Table S3: Gene signaling pathways related to GBM.
@@
rs1267622 (PDGFRA, PDGFB), (PDGRFA, SOS2), (GADD45A, GADD45B)
rs10519201 (E2F3, PIK3CA), (E2F3, BAX)
rs10488141 (SOS2, E2F1), (SOS2, GADD45B)
rs2076655 (E2F1, CALML5)
rs10509346 (E2F3, CALML5)
rs10518759 (E2F1, PIK3CA)
rs10512510 (E2F3, BRAF)
rs9303504 (E2F1, CALML5)
rs4834352 (E2F1, E2F3)
rs2075109 (PIK3CA, CALML5)
Table S4: Estimated SNP effects on gene co-expressions according to cspine.
@@
Our proof strategies for Theorems 1 and 2 follow that of Zhang and Li [2022], with modifications made to accommodate our concatenated design matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$. First we state a few lemmas.
Lemma 1 (Bellec et al. [2018] Lemma 1).
@@
Lemma 2 (Graybill and Marsaglia [1957] Theorem F).
@@
Lemma 3 (Laurent and Massart [2000] Lemma 1).
@@
Lemma 4 (Vershynin [2011] Proposition 5.16).
@@
Lemma 5 comes from Theorem 4.1 in Kuchibhotla and Chakrabortty [2022] applied to marginally sub-Gaussian random vectors.
Lemma 5 (Kuchibhotla and Chakrabortty [2022] Theorem 4.1).
Let $\boldsymbol{\mathbf{Z}}_{1},\dotsc,\boldsymbol{\mathbf{Z}}_{n}$ be independent random vectors in $\mathbb{R}^{p}$. Assume each element of $\boldsymbol{\mathbf{Z}}_{i}$ is sub-Gaussian with bounded sub-Gaussian norm for all $i\in[n]$. Let $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}=\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n$ and $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}}=\operatorname{\operatorname{\mathbb{E}}}[\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n]$. Define
@@
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}})\boldsymbol{\mathbf{v}}\right\rvert\precsim k\sqrt{\frac{\Upsilon_{n}\log p}{n}}+\frac{k\log n\log p}{n}$
with probability at least $1-O(1/p)$.
We may apply Lemma 5 because by Assumption 3, $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ is elementwise sub-Gaussian, each entry being the product of a sub-Gaussian and a bounded random variable. Furthermore, we have by the Cauchy-Schwarz inequality that
$\max_{\ell,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{i\ell}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{ik}\right)\leq\max_{\ell_{1},\ell_{2},\ell_{3},\ell_{4}}\operatorname{\operatorname{\mathbb{E}}}\left(X_{\ell_{1}}^{(1)^{2}}X_{\ell_{2}}^{(1)^{2}}U_{\ell_{3}}^{(1)^{2}}U_{\ell_{4}}^{(1)^{2}}\right)=O(1)$
@@
Define the matrices:
$\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}],\quad\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}}).$
@@
For a set of indices $S\subset[p(q+1)-1]$, denote by $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{S}$ the submatrix of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ with columns indexed by $S$. Let $\lVert\cdot\rVert_{\mathrm{op}}$ denote the matrix operator norm. Under Assumptions 1-4, there exist constants $M_{uw}$ and $C_{0}$ such that with probability at least $1-C_{0}\exp(-\log(pq))$ we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{S}\right\rVert_{\mathrm{op}}^{2}\leq M_{uw}$
@@
Letting $k=\lvert S\rvert$, it suffices to show that $\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ is bounded. We may write
@@
By Assumption 2, the second term is bounded. For the first term, by Lemma 5 we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\}\precsim\left(\frac{k^{2}\log(pq)}{n}\right)^{1/2}+\frac{k\log(pq)}{n/\log n}$
with probability at least $1-C_{0}\exp(-\log(pq))$ and we see that the right-hand side is bounded by Assumption 4. ∎
S4.1 Proof of Theorem 1
@@
In this step we bound the error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ by the stochastic term $\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle/n$, which is then bounded by a projection of $\boldsymbol{\mathbf{\varepsilon}}$ onto the columns of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$.
@@
Now let $\mathcal{I}$ and $\mathcal{J}$ be arbitrary index sets of the columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{W}}$ respectively. Denote by $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ the orthogonal projection onto the columns of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ indexed by $(\mathcal{I},\mathcal{J})$. Let $\mathcal{I}_{0}=S_{\gamma}\cup\hat{S}_{\gamma}$ and $\mathcal{J}_{0}=S_{\beta}\cup\hat{S}_{\beta}$ denote the unions of the true and estimated support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$. We seek to bound the stochastic term

### v2

@@ 6 GBM eQTL analysis
Among the SNPs identified as affecting the conditional dependence network, six are detected by both methods. We highlight two reprepresentative examples of these SNPs and their estimated effects on the network.
The first is rs10505980, a variant local to the KRAS gene, whose estimated effects are shown in Figure 5. Both methods detect that this variant influences the co-expression between PIK3CA and SHC2, genes involved in the Ras–Raf–MEK–ERK signaling pathway. In addition, cspine identifies an effect on the co-expression between PIK3CB and CAMK1G. Notably, PIK3CA and PIK3CB are key components of the PI3K/MTOR pathway, and their mutations are known to be oncogenic, with targeted inhibition being an active area of cancer research [samuels, whale_functional_2017]. In particular, inhibition of this pathway has been associated with improved survival in glioblastoma (GBM) [batsios_pi3kmtor_2019]. These findings suggest that rs10505980 may play a role in regulating the co-expression of oncogenic pathways involving PIK3CB and CAMK1G.
A second SNP identified by both methods is rs1267622, located near BRAF, with estimated effects shown in Figure 6. Both methods suggest that this variant modifies the co-expression between PDGFRA and SOS2. Additionally, RegGMM detects an effect on the co-expression between PDGFB and SOS2. Platelet-derived growth factors (PDGFA/B) and their receptors (PDGFRA/B) are known to play important roles in gliomas, where receptor activation triggers the Ras–Raf–MEK–ERK signaling cascade, in which SOS2 participates [soni_genetics_2005, westermark_platelet-derived_2014]. These results suggest that rs1267622 may influence the co-expression of genes in the PDGF receptor pathway and its downstream signaling processes.
Figure 5: Estimated covariate networks for the SNP rs10505980. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (9) and (3). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of PIK3CA and SHC2 is detected by both cspine and RegGMM. Our method cspine additionally identifies an effect on the co-expression of PIK3CB and CAMK1G.
Figure 6: Estimated covariate networks for the SNP rs1267622. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (3) and (9). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of GADD45A and GADD45B is detected by cspine while RegGMM detects an effect between PDGFRA and SOS2. The remaining two edges, between PDGRFA and PDGFB and PDGRFA and SOS2, are detected by both methods.
@@ 7 Discussion
In this work, we propose a new framework for covariate-adjusted Gaussian graphical models based on a natural parametrization that yields a jointly convex formulation. This approach enables unified estimation of both the covariated-dependent mean and covariate-dependent precision matrix, in contrast to the two-stage procedure of zhangli. Our theoretical analysis shows that the convex formulation permits more relaxed scaling conditions on the sparsity level of the true parameter relative to the sample size $n$. These advantages are supported by our simulation results, which demonstrate improved performance in regimes where stagewise methods are challenged by error propagation.
The proposed framework is flexible and does not impose restrictions on the form of the penalty functions $g_{j}$ in (13), allowing for a wide range of sparsity structures. Incorporating more structured penalties is a promising direction for future work. For instance, one may consider hierarchical penalties that reflect the relationship between covariate effects on the mean and precision matrix, where a covariate influencing the precision matrix is expected to also affect the mean, but not necessarily vice versa. Another direction is to develop tuning-free variants of the proposed method, for example by adapting ideas from the square-root lasso [sqrtlasso]. Additionally, methods for statistical inference could be developed by applying debiasing techniques to each nodewise regression [meng_statistical_2025]. From a theoretical perspective, it would be of interest to establish minimax optimality of the estimator in (14). While our results achieve the same rate as in zhangli, the optimality of this rate remains an open question.
@@ Acknowledgements
Data availability statement
The code and data supporting Section 5 can be found on the corresponding author’s GitHub. [^1] https://github.com/roobnloo/cspine-sim The data that support the findings Section 6 are available from the corresponding author, RL, upon reasonable request.
@@
Convex estimation of Gaussian graphical regression models with covariates
@@
$\boldsymbol{\mathbf{W}}_{-j}\in\mathbb{R}^{n\times(p-1)(q+1)}$ matrix of interactions.
@@
for $j=1,2,\dotsc,p$ do
   solve
@@
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\}}+\tilde{\beta}_{kjh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\}},\quad j\neq k$
@@
Table S3 displays the genes belonging to known signaling pathways associated with GBM. The table is copied from zhangli. Table S4 lists the SNPs that are detected by cspine to affect network co-expression.
@@
CALM1,CALML3, CALML4, CALML5, maklad
@@
Table S3: Gene signaling pathways related to GBM copied from zhangli and represented in https://www.genome.jp/pathway/hsa05214.
@@
cspine RegGMM
rs1267622 (SOS2, PDGFRA) (PDGFRA, PDGFB) (SOS2, PDGFB) (PDGFB, PDGFRA)
rs6701524 (AKT1, PIK3CD) (AKT1, MAP2K2) (MAP2K2, AKT1)
rs10505980 (PIK3CB, CAMK1G) (PIK3CA, SHC2) (PIK3CA, SHC2)
rs10519201 (E2F3, PIK3CA) (CDK6, PIK3CA)
rs10492796 (CDKN2A, CAMK2A) (CDKN2A, CAMK2A)
rs10488137 (CDKN2A, CAMK2A) (CDKN2A, CAMK2A)
rs10512510 (PIK3CD, PDGFRB) (E2F1, SOS1)
rs723211 (GADD45B, IGF1R)
rs723210 (GADD45B, AKT3)
rs3730504 (GADD45B, IGF1R)
rs198563 (CDKN2A, PIK3R3)
rs5871 (PIK3CB, CALML5)
rs10518759 (PIK3CA, CAMK1G)
rs1567810 (PIK3CB, IGF1R)
rs9303511 (SOS1, SHC2)
rs2078153 (PIK3R3, EGFR)
rs958686 (GADD45B, SOS1)
rs10496180 (GADD45B, CAMK2A)
rs559160 (CDKN2A, PIK3R3)
rs10488881 (GADD45B, AKT3)
rs759947 (PIK3CD, NRAS)
rs759948 (MAP2K1, PDGFRA)
rs727923 (PIK3R3, CAMK2A)
rs42670 (MAP2K1, CALML5)
rs39605 (DDB2, CALML5)
rs10488141 (PIK3CB, SOS2)
rs7785163 (CALML5, IGF1R)
rs1267600 (CALML5, PDGFRB)
rs10509346 (CALML5, PDGFRA)
rs473698 (CDK6, CAMK2A)
rs503314 (CCND1, RAF1)
rs10489985 (HRAS, SOS1)
rs2051785 (CDK6, AKT3)
rs9326835 (CDK6, CAMK2A)
Table S4: Estimated effects of SNPs on gene co-expressions.
@@
Our proof strategies for Theorems 1 and 2 follow that of zhangli, with modifications made to accommodate our concatenated design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$. We will first recall several lemmas from the literature and then prove some supporting lemmas before moving onto the main proofs.
S4.1 Technical lemmas
Lemma 1 (bellec Lemma 1).
@@
Lemma 2 (graybill Theorem F).
@@
Lemma 3 (laurent Lemma 1).
@@
Lemma 4 (vershynin Proposition 5.16).
@@
Lemma 5 comes from Theorem 4.1 in kuchi applied to marginally sub-exponential random vectors.
Lemma 5 (kuchi Theorem 4.1).
Let $\bm{Z}_{1},\dotsc,\bm{Z}_{n}$ be independent random vectors in $\mathbb{R}^{p}$ and let $\boldsymbol{\mathbf{Z}}=[\bm{Z}_{1},\dotsc,\bm{Z}_{n}]^{\top}$. Assume each element of $\bm{Z}_{i}$ is sub-exponential with bounded sub-exponential norm for all $i\in[n]$; $\lVert Z_{ij}\rVert_{\psi_{1}}=\sup_{d\geq 1}d^{-1}(\operatorname{\operatorname{\mathbb{E}}}\lvert Z_{ij}\rvert^{d})^{1/d}\leq K_{z}$ for all $i\in[n]$ and $j\in[p]$ for some constant $K_{z}>0$. Let $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}=\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n$ and $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}}=\operatorname{\operatorname{\mathbb{E}}}[\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n]$. Define
@@
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}})\boldsymbol{\mathbf{v}}\right\rvert\precsim k\sqrt{\frac{\Upsilon_{n}\log p}{n}}+K_{z}^{2}\frac{k(\log n\log p)^{2}}{n}$
with probability at least 1O(1/p)1-O(1/p).
With Assumptions 1 and 2, we may apply Lemma 5 to our design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ because it is elementwise sub-exponential, each entry being the product of a sub-Gaussian and a bounded random variable. Furthermore, we have by the Cauchy-Schwarz inequality that
$\max_{\ell,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{i\ell}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{ik}\right)\leq\max_{\ell_{1},\ell_{2},\ell_{3},\ell_{4}}\operatorname{\operatorname{\mathbb{E}}}\left(X_{\ell_{1}}^{(1)^{2}}X_{\ell_{2}}^{(1)^{2}}U_{\ell_{3}}^{(1)^{2}}U_{\ell_{4}}^{(1)^{2}}\right)=O(1)$
@@
Treating $j$ as fixed and suppressing the dependence of $\boldsymbol{\mathbf{W}}_{-j}$ on $j$, recall that the sample and population Gram matrices are defined as
$\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}],\quad\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left(\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\right).$
@@
For a set of indices $S\subset[p(q+1)-1]$, denote by $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{S}$ the submatrix of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ with columns indexed by $S$. Under Assumptions 1-4, there exist constants $M_{uw}$ and $C_{0}$ such that with probability at least $1-C_{0}\exp(-\log(pq))$ we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{S}\right\rVert_{\mathrm{op}}^{2}\leq M_{uw}$
@@
Letting $k=\lvert S\rvert$, it suffices to show that $\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ is bounded with the provided probability. We may write
@@
The second term is bounded by Assump 3. For the first term, by Lemma 5 we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\}\precsim k\left(\frac{\log(pq)}{n}\right)^{1/2}+\frac{k\log(pq)^{2}}{n/(\log n)^{2}}$
with probability at least $1-C_{0}\exp(-\log(pq))$ and we see that the right-hand side is $o(1)$ by Assumption 4. ∎
S4.2 Theorem 1 proof sketch
The proof of Theorem 1 closely follows that of Theorem 1 in zhangli. We give a sketch of the main proof technique, omitting constants while focusing on the important quantities. As such, the inequalities in this subsection should not be taken literally. Letting $\boldsymbol{\mathbf{\nu}}=\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\Delta}}=\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}$ denote the error terms, we first bound the sample squared error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ by the complexity term $E_{j}=(\sigma^{2}_{\varepsilon}/n)\cdot\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$. Recall the penalty term is $\mathrm{pen}=\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$. The proof occurs in three steps.
In the first step, a basic inequality is used to bound the sample squared error by the stochastic term plus the penalty term,
$\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\epsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\mathrm{pen}.$
The stochastic term is further bounded by the sample squared error and a projection $\mathcal{P}$ of error terms onto the support set of $\boldsymbol{\mathbf{\nu}}$ and $\boldsymbol{\mathbf{\Delta}}$:
$\langle\boldsymbol{\mathbf{\epsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle\leq\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lVert\mathcal{P}\boldsymbol{\mathbf{\epsilon}}\rVert_{2}^{2}.$
In the second step, by a counting argument and with $\chi^{2}$ tails, the projection term is shown to satisfy
$\lVert\mathcal{P}\boldsymbol{\mathbf{\epsilon}}\rVert_{2}^{2}\leq E_{j}+\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{r}{n}$
where $r$ is a high probability term with an exponentially small tail. Combining the above yields
$\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\leq E_{j}+\mathrm{pen}$
with high probability.
In the third step, the penalty term is shown to satisfy
$\mathrm{pen}\leq E_{j}+\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}.$
Lemmas 5 and 6 are then used to bound the difference between sample squared error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ and the population squared error $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$ by the estimation error. Combined with the above, this yields an upper bound on the population error as
$\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}(E_{j}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}).$
for a large constant $L$. Finally Assumption 3 is applied to get $\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\precsim E_{j}$ whence the desired bound.
S4.3 Proof of Theorem 1
@@
In this step we bound the error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ by the stochastic term $\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle/n$, which is then bounded by a projection of $\boldsymbol{\mathbf{\varepsilon}}$ onto the columns of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$.
@@
Now let $\mathcal{I}$ and $\mathcal{J}$ be arbitrary index sets of the columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{W}}$ respectively. Denote by $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ the orthogonal projection onto the columns of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ indexed by $(\mathcal{I},\mathcal{J})$. Let $\mathcal{I}_{0}=S_{\gamma}\cup\hat{S}_{\gamma}$ and $\mathcal{J}_{0}=S_{\beta}\cup\hat{S}_{\beta}$ denote the unions of the true and estimated support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$. We seek to bound the stochastic term